# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v2

@@ 3.2 Busemann points and 1-Lipschitz functions
It will be convenient to fix a linear form $p$ in the interior of the dual cone of $C$, and to consider the cross-section $C_{p}=\{x\in\operatorname{int}{C}\mid p(x)=1\}$. Then, $(x,y)\mapsto\operatorname{Hil}(x,y)$ is a bona fide metric on $C_{p}$. Observe that all the Hilbert balls of $C_{p}$ are compact, so that 34 can be applied. The group of isometries of $(C_{p},\operatorname{Hil})$ has been characterized in [LW11, Theorem 1.2] when $C=\mathbb{R}_{\geq}^{n}$ is the standard orthant, so that $C_{p}$ is the simplex. Let $\jmath$ denote the self-map of $(\mathbb{R}_{>0})^{n}$ such that $(\jmath(x))_{i}=x_{i}^{-1}$ for all $i\in[n]$, and $\rho$ denote the map from $\mathbb{R}_{>0}^{n}$ to $C_{p}$ such that $\rho(x)=x/p(x)$. Then, every isometry of $(C_{p},\operatorname{Hil})$ is either of the form $\phi(x)=\rho(DPx)$ where $D$ is a diagonal matrix with positive diagonal entries and $P$ is a permutation matrix, or of the form $\phi(x)=\rho(\jmath(DPx))$. Then, 34 shows that every (max,+)-isomorphism of the space of nonexpansive maps defined on $(C_{p},\operatorname{Hil})$ is of the form (30) where the function $g$ is constant. 34 also allows to characterize the set of (max,+)-isomorphisms of the space of real nonexpansive maps defined on $(C_{p},\operatorname{RFunk})$. When $n\geq 3$, considering $\delta=\delta^{\prime}=\operatorname{RFunk}$ in (30), it is straightforward to check that $g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))=\delta(x,x^{\prime})$ holds for all $x,x^{\prime}\in C_{p}$ iff $g$ is constant and $\phi$ is of the form $\phi(x)=\rho(DPx)$ (the case in which $\phi=\rho\circ\jmath(DPx)$ is excluded owing to the “unilateral” character of the Funk metric).
@@ 3.3 Continuous affine maps and the lower semicontinuous convex functions
We conclude studying the order isomorphisms over the set of lower semicontinuous convex functions, extending the result of [AAM09] to the generality of locally convex Hausdorff spaces. Our novel ingredient is that we identify in 41 the sup-irreducible points of the space of proper convex lower semicontinuous functions.
@@
denoting by $P^{f}\subset\mathscr{X}^{*}\times\mathbb{R}$ the nonempty set over which the supremum is taken. Fix $x_{0},x_{1}\in\mathscr{X}$, being possibly equal, and $(p_{0},\lambda_{0})\in P^{f}$, $(p_{1},\lambda_{1})\in P^{f}$. Assume that $p_{0}\neq p_{1}$, then by Hahn-Banach’s theorem applied to the locally convex space $\mathscr{X}^{*}$ equipped with the weak-* topology, we can take a weak-* closed halfspace $H^{+}$ containing $p_{0}$ but not $p_{1}$ and denote by $H^{-}$ the closure of its complement. This means that $H^{-}=\{p\in\mathscr{X}^{*}\,|\,\langle p,x_{H}\rangle\leq\alpha_{H}\}$ for some $x_{H}\in\mathscr{X}\backslash\{0\}$ and $\alpha_{H}\in\mathbb{R}$.
@@
Hence $\tilde{f}$ is a continuous affine function, and so is $f=\tilde{f}+\langle\bar{p},\cdot\rangle$. ∎
The following theorem has previously been shown for $\mathscr{X}=\mathbb{R}^{d}$ in [AAM09, Corollary 8]. Its generalization to Banach spaces appeared in [IRS15, Theorem 1]. All proofs rely eventually on a key result of affine geometry: in dimension greater than 2 only affine maps preserve segments. It should be highlighted that it is already stated in [AAM09, p670], before their Remark 9, that the characterization of anti-involutions given in 43 could be obtained from that of order isomorphisms as in 42. Artstein-Avidan and Milman tackle first anti-involutions and then the isomorphisms. As shown in 13, the difference between the two is that isomorphisms send relatively-inf-irreducible elements (here the indicator functions) onto themselves, while anti-involutions send them onto the sup-irreducible elements (here the continuous affine functions). This explains the main discrepancies of our proof with that of [AAM09, Theorem 5] and its similarity with that of [IRS15, Theorem 1]. To the difference of [IRS15], an argument based on weak-* continuity enables us to avoid all together considering the bidual in the proof. Aside from this result, our proof method is different because we emphasize the role of the sup-irreducible elements to obtain the characterization. Concerning anti-involutions, [CL21, Theorem 1.3] shows that reflexivity and a linear isomorphism between $\mathscr{X}$ and $\mathscr{X}^{*}$ are necessary and sufficient to have the Fenchel transform as anti-involution, as otherwise the bi-Fenchel is defined on the larger set $\mathscr{X}^{**}$.
@@
Using 13 and 41, we have that $J$ maps continuous affine functions onto continuous affine functions since they are the sup-irreducible points in this case. Let $G:\mathscr{X}^{*}\times\mathbb{R}\rightarrow\mathscr{X}^{*}\times\mathbb{R}$ be the bijective function that maps the coefficients $(p,\lambda)$ of an affine function $f$ onto the coefficients $(p^{\prime},\lambda^{\prime})$ of $Jf$. Let us write $G(p,\lambda)=(G_{1}(p,\lambda),G_{2}(p,\lambda))$
@@
with equality at some point $x_{0}$. Consequently, since $J$ commutes with the max operation by 7, we have that $Jf_{\alpha}\leq\max(Jf_{0},Jf_{1})$ with all functions being affine. Let us show that there exists a point $x^{\prime}_{0}$ where there is equality.
@@
Hence $d\geq 0$. Since $G$ is bijective in $\lambda$, we have that $d>0$ necessarily. It remains to study the continuity of the maps $A$ and $C$. Since $\mathcal{G}$ contains the $\{\delta_{x}^{\top}+\lambda\}_{x\in\mathscr{X},\lambda\in\overline{\mathbb{R}}}$, we derive from the intermediary result (29) in the proof of 30 that there exists a lower semicontinuous $g:\mathscr{X}\times\mathbb{R}\to\mathbb{R}$, with, for all $x\in\mathscr{X}$, $g(x,\cdot):\mathbb{R}\to\mathbb{R}$ continuous (being increasing and bijective), and a bijective $\phi:\mathscr{X}\to\mathscr{X}$ such that
@@
so, considering nets, we have that $A:\mathscr{X}^{*}\to\mathscr{X}^{*}$ is weak-* continuous as $g(x,\cdot)$ is continuous. Hence $A$ has an adjoint $A^{*}:\mathscr{X}\to\mathscr{X}$ with closed graph. Indeed, by weak-* continuity, for all $x\in\mathscr{X}$, there exists $y\in\mathscr{X}$ such that for all $p\in\mathscr{X}^{*}$, $\langle Ap,x\rangle=\langle p,y_{x}\rangle$. Set $A^{*}(x):=y_{x}$, since $\langle Ap,x\rangle=\langle p,A^{*}(x)\rangle$ and $\mathscr{X}^{*}$ separates the points of the locally convex $\mathscr{X}$, $A^{*}$ is linear. Take a net $(x_{\alpha},y_{\alpha})_{\alpha}$ with $y_{\alpha}=A^{*}x_{\alpha}$ converging to some $(x,y)\in\mathscr{X}\times\mathscr{X}$, then $\langle p,y\rangle=\langle Ap,x\rangle=\langle p,A^{*}x\rangle$, so $y=A^{*}x$ and the graph of $A^{*}$ is closed. Similarly if $(x_{\alpha})_{\alpha}$ converges weakly to some $x\in\mathscr{X}$, then $(A^{*}x_{\alpha})_{\alpha}$ converges weakly to $A^{*}x$.
Fix any convex l.s.c $f$. Then
$Jf(x)$ =(36)${{=}}\sup_{(p,\lambda)\in P^{f}}J(\langle p,\cdot\rangle+\lambda)(x)$ )
@@
Applying 11, we know that, for $\mathcal{G}=\operatorname*{Rg}(B)$ for some $\bar{B}$ as in (10), the anti-involutions are only the composition of the reference one $\bar{B}$ with order isomorphisms. In this case, the reference is the Fenchel transform, and we thus recover [AAM09, Theorem 5] for $\mathscr{X}=\mathbb{R}^{d}$ as a special case and derive a straightforward extension to semi-convex functions, i.e. the functions $f$ such that $f+\frac{\alpha}{2}\|\cdot\|^{2}$ is convex l.s.c.
@@
For convex functions, applying 11, since the Fenchel transform is an anti-involution with these assumptions on $\mathscr{X}$, see e.g. [CL21, Theorem 1.3], we have that $Tf(x)=\langle b,x\rangle+\delta+d\cdot f^{*}(K(Ax+c))$. The operator $TT$ is a (max,+)-isomorphism since $TT=\operatorname*{Id}_{\mathcal{G}}$. However $TT(f+\lambda)=TTf+d^{2}\lambda$, so necessarily $d=1$. Moreover $\{\langle p,\cdot\rangle+\lambda\}_{p\in\mathscr{X}^{*},\lambda\in\mathbb{R}}$ sup-generates $\mathcal{G}$, so $TT=\operatorname*{Id}_{\mathcal{G}}$ iff, for all $p\in\mathscr{X}^{*}$, $TT\langle p,\cdot\rangle=\langle p,\cdot\rangle$. Fix $p\in\mathscr{X}^{*}$, we have that $T\langle p,\cdot\rangle(x)=\langle b,x\rangle+\delta+\delta^{\top}_{A^{-1}(K^{-1}p-c)}$, and $TT\langle p,\cdot\rangle(x)=\langle b,x\rangle+\langle K(Ax+c)-b,A^{-1}(K^{-1}p-c)\rangle$. Identifying the “monomials” in $p$ and $x$, we get that $b=Kc$ and $K^{-1}A^{-\top}KA=\operatorname*{Id}_{\mathscr{X}}$ are necessary and sufficient conditions.
For the semi-convex functions, we just use 15 since by definition a function $f$ is $\alpha$-semiconvex iff $f+\frac{\alpha}{2}\|\cdot\|_{\mathscr{X}}^{2}$ is convex. ∎
@@ Acknowledgments.
The authors deeply thank Sebastián Tapia-García for the fruitful discussions and his many insights on topology. Pierre-Cyril Aubin-Frankowski was funded by the FWF project P 36344-N.
@@
Hence $\tilde{F}$ is continuous in the sense of 8 and commutes with the addition of $\overline{\mathbb{R}}$-scalars. We can now apply 9 to derive a kernel $b$ associated with $\tilde{F}$. Since $\tilde{F}^{1}\bar{F}=\operatorname*{Id}_{\mathscr{F}}$, we have that $\mathscr{F}=\tilde{F}^{1}\operatorname*{Rg}(\tilde{F})$, so $\mathcal{G}=\operatorname*{Rg}(\tilde{F})$. As $\bar{F}$ is an anti-isomorphism,
@@
We have thus recovered one of the characterizations of dual Galois connections over ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$, as given in [AGK05, p.3, Eq.(2a)]. [AGK05, Theorem 2.1, Example 2.8] then allows to conclude, since the dual connection is simply the transpose $\tilde{F}^{\top}$ [Sin97, Theorem 8.4].

### v3

@@ 3.2 Busemann points and 1-Lipschitz functions
It will be convenient to fix a linear form $p$ in the interior of the dual cone of $C$, and to consider the cross-section $C_{p}=\{x\in\operatorname{int}{C}\mid p(x)=1\}$. Then, $(x,y)\mapsto\operatorname{Hil}(x,y)$ is a bona fide metric on $C_{p}$. Observe that all the Hilbert balls of $C_{p}$ are compact, so that ˜33 can be applied. The group of isometries of $(C_{p},\operatorname{Hil})$ has been characterized in [LW11, Theorem 1.2] when $C=\mathbb{R}_{\geq}^{n}$ is the standard orthant, so that $C_{p}$ is the simplex. Let $\jmath$ denote the self-map of $(\mathbb{R}_{>0})^{n}$ such that $(\jmath(x))_{i}=x_{i}^{-1}$ for all $i\in[n]$, and $\rho$ denote the map from $\mathbb{R}_{>0}^{n}$ to $C_{p}$ such that $\rho(x)=x/p(x)$. Then, every isometry of $(C_{p},\operatorname{Hil})$ is either of the form $\phi(x)=\rho(DPx)$ where $D$ is a diagonal matrix with positive diagonal entries and $P$ is a permutation matrix, or of the form $\phi(x)=\rho(\jmath(DPx))$. Then, ˜33 shows that every (max,+)-isomorphism of the space of nonexpansive maps defined on $(C_{p},\operatorname{Hil})$ is of the form (29) where the function $g$ is constant. ˜33 also allows to characterize the set of (max,+)-isomorphisms of the space of real nonexpansive maps defined on $(C_{p},\operatorname{RFunk})$. When $n\geq 3$, considering $\delta=\delta^{\prime}=\operatorname{RFunk}$ in (29), it is straightforward to check that $g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))=\delta(x,x^{\prime})$ holds for all $x,x^{\prime}\in C_{p}$ iff $g$ is constant and $\phi$ is of the form $\phi(x)=\rho(DPx)$ (the case in which $\phi=\rho\circ\jmath(DPx)$ is excluded owing to the “unilateral” character of the Funk metric).
@@ 3.3 Continuous affine maps and the lower semicontinuous convex functions
We conclude by studying the order isomorphisms over the set of lower semicontinuous convex functions, extending the result of [AAM09] to the generality of locally convex Hausdorff spaces. Our novel ingredient is that we identify in ˜40 the sup-irreducible points of the space of proper convex lower semicontinuous functions.
@@
[…]
@@
Consequently $\tilde{f}$ is a continuous affine function, and so is $f=\tilde{f}+\langle\bar{p},\cdot\rangle$. ∎
The following theorem has previously been shown for $\mathscr{X}=\mathbb{R}^{d}$ in [AAM09, Corollary 8]. Its generalization to Banach spaces appeared in [IRS15, Theorem 1]. All proofs rely eventually on a key result of affine geometry: in dimension greater than 2 only affine maps preserve segments. It should be highlighted that it is already stated in [AAM09, p670], before their Remark 9, that the characterization of anti-involutions given in ˜42 could be obtained from that of order isomorphisms as in ˜41. Artstein-Avidan and Milman tackle first anti-involutions and then the isomorphisms. As shown in ˜13, the difference between the two is that isomorphisms send relatively-inf-irreducible elements (here the indicator functions) onto themselves, while anti-involutions send them onto the sup-irreducible elements (here the continuous affine functions). This explains the main discrepancies of our proof with that of [AAM09, Theorem 5] and its similarity with that of [IRS15, Theorem 1]. In constrast to [IRS15], an argument based on weak-* continuity enables us to avoid all together considering the bidual in the proof. Aside from this result, our proof method is different because we emphasize the role of the sup-irreducible elements to obtain the characterization. Concerning anti-involutions, [CL21, Theorem 1.3] shows that reflexivity and a linear isomorphism between $\mathscr{X}$ and $\mathscr{X}^{*}$ are necessary and sufficient to have the Fenchel transform as anti-involution, as otherwise the bi-Fenchel is defined on the larger set $\mathscr{X}^{**}$. Using different proof techniques specific to the convex case, an even more general result to ours was obtained recently in [Leu23, Theorem 2.13] with points in convex subsets of locally convex Hausdorff space and values in ordered topological vector spaces.
@@
Using ˜13 and ˜40, we have that $J$ maps continuous affine functions onto continuous affine functions since they are the sup-irreducible points in this case. Let $G:\mathscr{X}^{*}\times\mathbb{R}\rightarrow\mathscr{X}^{*}\times\mathbb{R}$ be the bijective function that maps the coefficients $(p,\lambda)$ of an affine function $f$ onto the coefficients $(p^{\prime},\lambda^{\prime})$ of $Jf$. Let us write $G(p,\lambda)=(G_{1}(p,\lambda),G_{2}(p,\lambda))$
@@
with equality at some point $x_{0}$. Consequently, since $J$ commutes with the max operation by ˜7, we have that $Jf_{\alpha}\leq\max(Jf_{0},Jf_{1})$ with all functions being affine. Let us show that there exists a point $x^{\prime}_{0}$ where there is equality.
@@
Hence $d\geq 0$. Since $G$ is bijective in $\lambda$, we have that $d>0$ necessarily. It remains to study the continuity of the maps $A$ and $C$. Since $\mathcal{G}$ contains the $\{\delta_{x}^{\top}+\lambda\}_{x\in\mathscr{X},\lambda\in\overline{\mathbb{R}}}$, we derive from the intermediary result (28) in the proof of ˜29 that there exists a lower semicontinuous $g:\mathscr{X}\times\mathbb{R}\to\mathbb{R}$, with, for all $x\in\mathscr{X}$, $g(x,\cdot):\mathbb{R}\to\mathbb{R}$ continuous (being increasing and bijective), and a bijective $\phi:\mathscr{X}\to\mathscr{X}$ such that
@@
so, considering nets, we have that $A:\mathscr{X}^{*}\to\mathscr{X}^{*}$ is weak-* continuous as $g(x,\cdot)$ is continuous. Hence $A$ has an adjoint $A^{*}:\mathscr{X}\to\mathscr{X}$ with closed graph. Indeed, by weak-* continuity, for all $x\in\mathscr{X}$, there exists $y_{x}\in\mathscr{X}$ such that for all $p\in\mathscr{X}^{*}$, $\langle Ap,x\rangle=\langle p,y_{x}\rangle$. Set $A^{*}(x):=y_{x}$, since $\langle Ap,x\rangle=\langle p,A^{*}(x)\rangle$ and $\mathscr{X}^{*}$ separates the points of the locally convex $\mathscr{X}$, $A^{*}$ is linear. Take a net $(x_{\alpha},y_{\alpha})_{\alpha}$ with $y_{\alpha}=A^{*}x_{\alpha}$ converging to some $(x,y)\in\mathscr{X}\times\mathscr{X}$, then $\langle p,y\rangle=\langle Ap,x\rangle=\langle p,A^{*}x\rangle$, so $y=A^{*}x$ and the graph of $A^{*}$ is closed. Similarly if $(x_{\alpha})_{\alpha}$ converges weakly to some $x\in\mathscr{X}$, then $(A^{*}x_{\alpha})_{\alpha}$ converges weakly to $A^{*}x$.
Fix any convex l.s.c. $f$. Then
[…]
@@
Applying ˜11, we know that, for $\mathcal{G}=\operatorname*{Rg}(B)$ for some $\bar{B}$ as in (10), the anti-involutions are of the form $\bar{B}J$ with order isomorphisms $J:\mathcal{G}\to\mathcal{G}$ such that $\bar{B}J\bar{B}J=\operatorname*{Id}_{\mathcal{G}}$. We can equivalently consider $J\bar{B}$. In the present case, $\bar{B}$ is the Fenchel transform, and we thus recover [AAM09, Theorem 5] for $\mathscr{X}=\mathbb{R}^{d}$ as a special case and derive a straightforward extension to semi-convex functions, i.e. the functions $f$ such that $f+\frac{\alpha}{2}\|\cdot\|^{2}$ is convex l.s.c.
@@
For convex functions, applying ˜11, since the Fenchel transform is an anti-involution with these assumptions on $\mathscr{X}$, see e.g. [CL21, Theorem 1.3], we have that $Tf(x)=\langle b,x\rangle+\delta+d\cdot f^{*}(K(Ax+c))$. The operator $TT$ is a (max,+)-isomorphism since $TT=\operatorname*{Id}_{\mathcal{G}}$. However $TT(f+\lambda)=TTf+d^{2}\lambda$, so necessarily $d=1$. Moreover $\{\langle p,\cdot\rangle+\lambda\}_{p\in\mathscr{X}^{*},\lambda\in\mathbb{R}}$ sup-generates $\mathcal{G}$, so $TT=\operatorname*{Id}_{\mathcal{G}}$ iff, for all $p\in\mathscr{X}^{*}$, $TT\langle p,\cdot\rangle=\langle p,\cdot\rangle$. Fix $p\in\mathscr{X}^{*}$, we have that $T\langle p,\cdot\rangle(x)=\langle b,x\rangle+\delta+\delta^{\top}_{A^{-1}(K^{-1}p-c)}$, and $TT\langle p,\cdot\rangle(x)=\langle b,x\rangle+\langle K(Ax+c)-b,A^{-1}(K^{-1}p-c)\rangle$. Identifying the “monomials” in $p$ and $x$, we get that $b=Kc$ and $K^{-1}A^{-\top}KA=\operatorname*{Id}_{\mathscr{X}}$ are necessary and sufficient conditions.
For the semi-convex functions, we just use ˜15 since by definition a function $f$ is $\alpha$-semiconvex iff $f+\frac{\alpha}{2}\|\cdot\|_{\mathscr{X}}^{2}$ is convex. ∎
@@ Acknowledgments.
The authors deeply thank Sebastián Tapia-García for the fruitful discussions and his many insights on topology; as well as a very thorough anonymous reviewer for their detailed reading which improved the paper. Most of the article was written while Pierre-Cyril Aubin-Frankowski was hosted at TU Wien, being funded by the FWF project P 36344-N.
@@
[…]
@@
We have thus recovered one of the characterizations of dual Galois connections over $\overline{\mathbb{R}}^{\mathscr{Y}}$, as given in [AGK05, p.3, Eq.(2a)]. [AGK05, Theorem 2.1, Example 2.8] then allows to conclude, since the dual connection is simply the transpose $\tilde{F}^{\top}$ [Sin97, Theorem 8.4].